# Distribution of lengths of closed saddle connections on moduli space of large genus translation surface

### v1

@@ Abstract.
Let $S_{g}$ be a closed surface of genus $g$ and ${\mathcal{H}}_{g}$ be the moduli space of Abelian differentials on $S_{g}$. A stratum of ${\mathcal{H}}_{g}$, together with the Masur-Veech measure, becomes a probability space. Then the number of closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ on a random flat surface in the stratum is a random variable. We prove that when $g\to\infty$, the distribution of the random variable converges to a Poisson distributed random variable. This result answers a question of Masur, Rafi and Randecker.
Keywords: Random flat surface; closed saddle connection
@@ 1. Introduction
Let $L_{g,[a,b]}(X,\omega)$ be the number of closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Denote by $\mathcal{C}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ consisting of flat surfaces having closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. The main result of this paper is the following theorem
@@ 2.1.1. Flat metric
For $(X,\omega)\in{\mathcal{H}}_{g}$, it defines a conical metric with singularities on $X$ given by
@@
A saddle connection on $(X,\omega)$ is an $|\omega|$- geodesic segment connecting two zeros of $\omega$ and has no other zeros in the interior. We call a saddle connection open if it connects two different zeros and closed if it connects a zero to itself. Generally, a geodesic on $(X,\omega)$ is a concatenation of some saddle connections, with the angle between two adjacent saddle connections larger than $\pi$. However, there are geodesics with no zeros on them, which we call regular.
@@ 2.1.2. Stratum and Masur-Veech measure
On each stratum, the period mapping induces a natural Lebesgue measure such that the lattice $H_{1}(X,\Sigma,\mathbb{Z}\oplus\mathbb{Z}i)$ has co-volume $1$, which is introduced by Masur [10] and Veech [16], we call it Masur-Veech measure. For a measured set $K$ of ${\mathcal{H}}_{g}(\kappa)$, denote its Masur-Veech measure by $\operatorname{Vol}(K)$.
Now we consider the hypersurface ${\mathcal{H}}^{1}_{g}(\kappa)$ of ${\mathcal{H}}_{g}(\kappa)$ consisting of $(X,\omega)$ with unit-area. The Masur-Veech measure on ${\mathcal{H}}^{1}_{g}(\kappa)$ can be defined as follows: for an open set $K$ on ${\mathcal{H}}^{1}_{g}(\kappa)$, its Masur-Veech measure is defined by
@@ 2.2.1. Configuration
For a flat surface $(X,\omega)$ and a saddle connection $\gamma$ on it, define the holonomy vector of $\gamma$ to be $\int_{\gamma}\omega$ and denote it by $hol(\gamma)$. Denote the set of all saddle connections of $(X,\omega)$ by $\tilde{V}_{sc}(X,\omega)$. Then define
@@
By [18, Proposition 3.1], $V_{sc}(X,\omega)$ is a discrete subset of $\mathbb{R}^{2}$ for any $(X,\omega)$. Note that different saddle connections may have same holonomy vector. For example two homologous saddle connections must have same holonomy, here homologous for two open sadddle connections $\beta$ and $\gamma$ means $\beta\circ\gamma^{-1}$ is homologous to zero. For a vector $\vec{v}\in V_{sc}(X,\omega)$, define the multiplicity of $\vec{v}$ to be $|hol^{-1}(\vec{v})|$. It is known that almost any flat surface in any stratum does not have a pair of non-homologous sharing a same holonomy vector by [7, Proposition 3.1]. So in the following we consider only the surfaces on which if two saddle connections have same holonomy, then they are homologous.
@@
For open saddle connections, first fix two zeros $p_{1}$ and $p_{2}$ with orders $m_{1}$ and $m_{2}$, consider the saddle connections connecting $p_{1}$ and $p_{2}$ of multiplicity $p$ and denote them by $\gamma_{1},\cdots,\gamma_{p}$, where the orientation is from $p_{1}$ to $p_{2}$ and the cyclic order of $\gamma_{i}$ is clockwise at $p_{1}$. Now we consider the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ and $p_{2}$, which are even multiples of $\pi$. Let the the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ be $2(a^{\prime}_{i}+1)\pi$ , the the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{2}$ be $2(a^{\prime\prime}_{i}+1)\pi$. Then a configuration $\mathcal{C}$ of open saddle connections consists of $(m_{1},m_{2},a^{\prime}_{i},a^{\prime\prime}_{i})$.
@@ 2.3. Factorial moment
$\mathbb{P}(X=k)=\frac{\lambda^{k}e^{\lambda}}{k!}.$
@@ 3. Some claims
Before the proof of the main theorem, we have to make some claim which will simplify the situation. We consider the stratum whose zero order is $O(1)$, and it inplies $|\kappa|=O(g)$.
@@
For a closed connection $\gamma$ of $(X,\omega)$, if it’s separable, then it cut $X$ into $X_{1}$ and $X_{2}$, and $\partial X_{1}=\partial X_{2}=\gamma$. By Stokes theorem, $\int_{\gamma}\omega=0$, but since $\gamma$ is a saddle connection its length is $|\int_{\gamma}\omega|$, which is a contradiction. ∎
@@
From Theorem 2.4, the Siegel-Veech constant of configuration with closed saddle connection involving $q$ cylinders of multiplicity $p$ is $O(\frac{1}{g^{2p+q-2}})$ . Denote by $\mathcal{C}^{p,q}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ which has a closed saddle connection of length in the interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ involving $q$ cylinders of multiplicity $p$. Then for $p\geq 2$ or $q\geq 1$, we have
@@
Combine claim $2,3$ and $4$, we have that for a closed saddle connection $\gamma$ on $(X,\omega)$ in principal stratum, since its total angle around the zero is $4\pi$ and has to be divided into odd multiples of $\pi$, it must have angle $\pi$ at one side of $\gamma$ and $3\pi$ at another. So there is a cyliner with $\gamma$ as a boundary. As for the other boundary, if it consists of some open saddle connections, then these saddle connections have angles $\pi$. That means there are at least two non-homologous saddle connections have angle $\pi$. The measure of the set of such surfaces is zero. So we only need to consider the cylinder bounded by curves homologous to $\gamma$. Then from Claim 2 the probability measure of $\mathcal{C}_{g}(1^{2g-2})$ has limit
@@
So we need consider the question on stratum with higher order zeros. Next we will consider ${\mathcal{H}}_{g}(2^{g-1})$ firstly.
@@ 4.1. Open up a higher-order zero
Cut along $\sigma+\sigma^{\prime}$ and denote the two copies of $\sigma+\sigma^{\prime}$ and $v_{1}$ by $\sigma^{\pm}+(\sigma^{\prime})^{\pm}$ and $v_{1}^{\pm}$. Then glue $\sigma^{+}$ and $(\sigma^{\prime})^{+}$, $\sigma^{-}$ and $(\sigma^{\prime})^{-}$. This surgery reduces the order of $v_{1}$ and constructs a new Abelian differential $(X^{\prime},\omega^{\prime})$. On $(X^{\prime},\omega^{\prime})$, the total angle at $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $2k_{1}\pi$, $2k_{2}\pi$, and $2(n_{2}+2)\pi$ respectively, that is, the order of $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $k_{1}-1$, $k_{2}-1$, and $n_{2}+1$. Particularly, if $v_{1}$ is a simple zero, the surgery collapses the saddle connection. We call this surgery collapsing surgery.
@@
In our situation, we need to use the inverse of collapsing surgery which we call it opening surgery. From Claim 2, it suffices to consider the subset $\mathcal{C}^{1,0}_{g}(2^{g-1})$. For an Abelian differential $(X,\omega)$ in $\mathcal{C}^{1,0}_{g}(2^{g-1})$, choose a closed saddle connection $\gamma$ at a double zero $p$. We can reverse the surgery introduced above as follows.
@@
This surgery is the inverse of the collapsing surgery above : the double zero is replaced by two simple zeros, and it can be carried out except for the surgery locus goes through some zero, that is, it has two non-homologous saddle connections with angle $2\pi$. From period mapping, such subset has measure zero. We call the subset on which the surgery can be carried out permissible set, and denote it by $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$, from the discussion above we have
@@
$\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$ is a full measure subset of $\mathcal{C}^{1,0}_{g}(2^{g-1})$.
@@ 4.2. Move zero along closed curve and pinch
$F_{2}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\to{\mathcal{H}}_{g}(1^{2},2^{g-1})$
@@ 4.2.2. $F_{2}$ is well-defined almost everywhere
We have defined a mapping if the saddle connections $\gamma_{1}$ and $\gamma_{2}$ are preserved along the curve $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$ on $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$. Next we will see the mapping can be defined for almost every flat surface in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
From above we know the moving surgery can be realized except for some of $\gamma_{1}$ or $\gamma_{2}$ degenerate along $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Suppose $T$ is the first time when the geodesic in $\gamma_{1}]$ can be represented by $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$, since the saddle connection is smooth on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$, the corner of $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$ must be $\pi$. But the holonomy of $pp_{i}$ is not changed on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$ This implies there exists some $pp_{i}$ such that its holonomy on $(X^{\prime},\omega^{\prime})$ has the same direction with $p_{0}p_{1}$. And by the relation of holonomy under $F_{1}$, this implies $p_{0}p_{i}$ has the same direction with the closed curve on $(X,\omega)$. Under period mapping, on each local chart this is a measure-zero set. So all such $(X,\omega)$ is a subset of measure zero in ${\mathcal{H}}_{g}(2^{g-1})$, also in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
@@ 4.2.3. Pinching
On the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$ where $F_{2}$ is well-defined, we can cut along $\gamma_{1}+\gamma_{2}$ and denote the two copies of $p$ and $q$ by $^{+}, qq^{-}$, $^{+}, pp^{-}$, the two copies of $\gamma_{1}$ and $\gamma_{2}$ by $\gamma^{+}_{1}$, $\gamma^{-}_{1}$, $\gamma^{+}_{2}$, $\gamma^{-}_{2}$. Notice that the total angles at $^{+}, qq^{-}$, $^{+}, pp^{-}$ are all $2\pi$, which means $\gamma^{\pm}_{1}$ have same direction with $\gamma^{\pm}_{2}$, then glue $\gamma^{+}_{1}$ with $\gamma^{+}_{2}$, $\gamma^{-}_{1}$ with $\gamma^{-}_{2}$, which make $^{+}, qq^{-}$, $^{+}, pp^{-}$ become regular points. Since $\gamma_{1}+\gamma_{2}$ is non-separable, this surgery reduces the genus by one, and obtains a new abelian differential on which we mark two points $p^{+}$ and $p^{-}$ so as to find its inverse, then we obtain:
$F_{3}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\to{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$
@@
which is a measure zero subset of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$.
@@
whose Jacobian has deteminant one and thus is measure-preserving. So we have
$F_{3}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1}))=[{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})]\setminus\tilde{{\mathcal{H}}}$,
Since $\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2},0,0))=\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2}))$, we have

### v2

@@ Abstract.
Let $S_{g}$ be a closed surface of genus $g$ and ${\mathcal{H}}_{g}$ be the moduli space of Abelian differentials on $S_{g}$. A stratum of ${\mathcal{H}}_{g}$, endowed with the Masur-Veech measure, becomes a probability space. Then the number of closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ on a random translation surface in the stratum is a random variable. We prove that when $g\to\infty$, the distribution of the random variable converges to a Poisson distributed random variable. This result answers a question of Masur, Rafi and Randecker.
Keywords: Random translation surface; closed saddle connection
@@ 1. Introduction
Let $L_{g,[a,b]}(X,\omega)$ be the number of closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Denote by $\mathcal{C}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ consisting of translation surfaces on which there exist closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. The main result of this paper is the following theorem
@@ 2.1.1. Flat metric
A surace $(X,\omega)\in{\mathcal{H}}_{g}$ is endowed with conical metric with singularities on $X$ given by
@@
A saddle connection on $(X,\omega)$ is an $|\omega|$- geodesic segment connecting two zeros of $\omega$ and has no other zeros in the interior. We call a saddle connection open if it connects two different zeros and closed if it connects a zero to itself. Generally, a geodesic on $(X,\omega)$ is a concatenation of some saddle connections, with the angle between two adjacent saddle connections larger than $\pi$. However, there are geodesics that do not pass through any zeros, which we call regular.
@@ 2.1.2. Stratum and Masur-Veech measure
On each stratum, the period mapping induces a natural Lebesgue measure, normalized so that the lattice $H_{1}(X,\Sigma,\mathbb{Z}\oplus\mathbb{Z}i)$ has co-volume $1$. This measure, which is introduced by Masur [8] and Veech [13], is called Masur-Veech measure. For a measured set $K$ of ${\mathcal{H}}_{g}(\kappa)$, denote its Masur-Veech measure by $\operatorname{Vol}(K)$.
Now we consider the hypersurface ${\mathcal{H}}^{1}_{g}(\kappa)$ of ${\mathcal{H}}_{g}(\kappa)$ consisting of $(X,\omega)$ with unit area. The Masur-Veech measure on ${\mathcal{H}}^{1}_{g}(\kappa)$ can be defined as follows: for an open set $K$ on ${\mathcal{H}}^{1}_{g}(\kappa)$, its Masur-Veech measure is defined by
@@ 2.2.1. Configuration
For a translation surface $(X,\omega)$ and a saddle connection $\gamma$ on it, define the holonomy vector of $\gamma$ to be $\int_{\gamma}\omega$ and denote it by $hol(\gamma)$. Denote the set of all saddle connections of $(X,\omega)$ by $\tilde{V}_{sc}(X,\omega)$. Then define
@@
By [15, Proposition 3.1], $V_{sc}(X,\omega)$ is a discrete subset of $\mathbb{R}^{2}$ for any $(X,\omega)$. Note that different saddle connections may have same holonomy vector. For example two homologous saddle connections must have same holonomy, here homologous for two open sadddle connections $\beta$ and $\gamma$ means $\beta\circ\gamma^{-1}$ is homologous to zero. For a vector $\vec{v}\in V_{sc}(X,\omega)$, define the multiplicity of $\vec{v}$ to be $|hol^{-1}(\vec{v})|$. It is known that almost any translation surface in any stratum does not have a pair of non-homologous sharing a same holonomy vector by [6, Proposition 3.1]. So in the following we consider only the surfaces on which if two saddle connections have same holonomy, then they are homologous.
@@
For open saddle connections, first fix two zeros $p_{1}$ and $p_{2}$ with orders $m_{1}$ and $m_{2}$, consider the saddle connections connecting $p_{1}$ and $p_{2}$ of multiplicity $p$ and denote them by $\gamma_{1},\cdots,\gamma_{p}$, where the orientation is from $p_{1}$ to $p_{2}$ and the cyclic order of $\gamma_{i}$ is clockwise at $p_{1}$. Now we consider the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ and $p_{2}$, which are even multiples of $\pi$. Let the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ be $2(a^{\prime}_{i}+1)\pi$ , the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{2}$ be $2(a^{\prime\prime}_{i}+1)\pi$. Then a configuration $\mathcal{C}$ of open saddle connections consists of $(m_{1},m_{2},a^{\prime}_{i},a^{\prime\prime}_{i})$.
@@ 2.3. Factorial moment
$\mathbb{P}(X=k)=\frac{\lambda^{k}e^{-\lambda}}{k!}.$
@@ 3. Some claims
Before the proof of the main theorem, we have to make some claims which will simplify the situation. We consider the stratum whose zero order is $O(1)$, and it inplies $|\kappa|=O(g)$.
@@
For a closed connection $\gamma$ of $(X,\omega)$, if it’s separable, then it cuts $X$ into $X_{1}$ and $X_{2}$, and $\partial X_{1}=\partial X_{2}=\gamma$. By Stokes theorem, $\int_{\gamma}\omega=0$, but since $\gamma$ is a saddle connection its length is $|\int_{\gamma}\omega|$, which is a contradiction. ∎
@@
From Theorem 2.4, the Siegel-Veech constant of configuration with closed saddle connection involving $q$ cylinders of multiplicity $p$ is $O(\frac{1}{g^{2p+q-2}})$ . Denote by $\mathcal{C}^{p,q}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ consisting of surfaces which have a closed saddle connection of length in the interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ involving $q$ cylinders of multiplicity $p$. Then for $p\geq 2$ or $q\geq 1$, we have
@@
Combine claim $2,3$ and $4$, we have that for a closed saddle connection $\gamma$ on $(X,\omega)$ in principal stratum, since its total angle around the zero is $4\pi$ and has to be divided into odd multiples of $\pi$, it must have angle $\pi$ at one side of $\gamma$ and $3\pi$ at another. So there is a cyliner with $\gamma$ as a boundary. As for the other boundary, if it consists of some open saddle connections, then these saddle connections have angles $\pi$. That means there are at least two non-homologous saddle connections have angle $\pi$, which occurs on a set of measure zero. So we only need to consider the cylinder bounded by curves homologous to $\gamma$. Then from Claim 2 the probability measure of $\mathcal{C}_{g}(1^{2g-2})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ has limit
@@
So we need to consider the question on stratum with higher order zeros. Next we will consider ${\mathcal{H}}_{g}(2^{g-1})$ firstly.
@@ 4.1. Open up a higher-order zero
Cut along $\sigma+\sigma^{\prime}$ and denote the two copies of $\sigma+\sigma^{\prime}$ and $v_{1}$ by $\sigma^{\pm}+(\sigma^{\prime})^{\pm}$ and $v_{1}^{\pm}$. Then glue $\sigma^{+}$ and $(\sigma^{\prime})^{+}$, $\sigma^{-}$ and $(\sigma^{\prime})^{-}$. This surgery reduces the order of $v_{1}$ and constructs a new Abelian differential $(X^{\prime},\omega^{\prime})$. On $(X^{\prime},\omega^{\prime})$, the total angles at $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $2k_{1}\pi$, $2k_{2}\pi$, and $2(n_{2}+2)\pi$ respectively, that is, the orders of $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $k_{1}-1$, $k_{2}-1$, and $n_{2}+1$. Particularly, if $v_{1}$ is a simple zero, the surgery collapses the saddle connection. We call this surgery collapsing surgery.
@@
In our situation, we need to use the inverse of collapsing surgery which we call opening surgery. From Claim 2, it suffices to consider the subset $\mathcal{C}^{1,0}_{g}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. For an Abelian differential $(X,\omega)$ in $\mathcal{C}^{1,0}_{g}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, choose a closed saddle connection $\gamma$ with length in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ at a double zero $p$. We can reverse the surgery introduced above as follows.
@@
This surgery is the inverse of the collapsing surgery above: the double zero is replaced by two simple zeros, and it can be carried out except for the surgery locus goes through some zero, that is, it has two non-homologous saddle connections with angle $2\pi$. From period mapping, such subset has measure zero. We call the subset on which the surgery can be carried out permissible set, and denote it by $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, from the discussion above we have
@@
$\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ is a full measure subset of $\mathcal{C}^{1,0}_{g}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$.
@@ 4.2. Move zero along closed curve and pinch
$F_{2}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\to{\mathcal{H}}_{g}(1^{2},2^{g-1})$
@@ 4.2.2. $F_{2}$ is well-defined almost everywhere
We have defined a mapping if the saddle connections $\gamma_{1}$ and $\gamma_{2}$ are preserved along the curve $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$ on $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$. Next we will see the mapping can be defined for almost every translation surface in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
From above we know the moving surgery can be realized except for some of $\gamma_{1}$ or $\gamma_{2}$ degenerate along $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Suppose $T$ is the first time when the geodesic in $\gamma_{1}]$ can be represented by $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$, since the saddle connection is smooth on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$, the corner of $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$ must be $\pi$. But the holonomy of $pp_{i}$ is not changed on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$ This implies there exists some $pp_{i}$ such that its holonomy on $(X^{\prime},\omega^{\prime})$ has the same direction with $p_{0}p_{1}$. And by the relation of holonomy under $F_{1}$, this implies $p_{0}p_{i}$ has the same direction with the closed curve on $(X,\omega)$. Under period mapping, on each local chart this is a measure-zero set. So all such $(X,\omega)$ is a subset of measure zero in ${\mathcal{H}}_{g}(2^{g-1})$, also in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$.
@@ 4.2.3. Pinching
On the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ where $F_{2}$ is well-defined, we can cut along $\gamma_{1}+\gamma_{2}$ and denote the two copies of $p$ and $q$ by $^{+}, qq^{-}$, $^{+}, pp^{-}$, the two copies of $\gamma_{1}$ and $\gamma_{2}$ by $\gamma^{+}_{1}$, $\gamma^{-}_{1}$, $\gamma^{+}_{2}$, $\gamma^{-}_{2}$. Notice that the total angles at $^{+}, qq^{-}$, $^{+}, pp^{-}$ are all $2\pi$, which means $\gamma^{\pm}_{1}$ have same direction with $\gamma^{\pm}_{2}$, then glue $\gamma^{+}_{1}$ with $\gamma^{+}_{2}$, $\gamma^{-}_{1}$ with $\gamma^{-}_{2}$, which make $^{+}, qq^{-}$, $^{+}, pp^{-}$ become regular points. Since $\gamma_{1}+\gamma_{2}$ is non-separable, this surgery reduces the genus by one, and obtains a new abelian differential on which we mark two points $p^{+}$ and $p^{-}$ so as to find its inverse. Considering the choice of which zero is selected after applying the inverse map, we finally obtain:
$F_{3}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\to M\times{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]},$
where $M$ is the combinatorial data of choosing one from g1g-1 zeros.
@@
which is a measure zero subset of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$. So we have
$F_{3}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1}))\subset[{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})]\setminus\tilde{{\mathcal{H}}}$.
@@
whose Jacobian has deteminant one and thus is measure-preserving.
5. Collapse closed saddle connections simultaneously
From the construction above we can collapse one closed saddle connection on some $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$ and obtain a new surface $(X^{\prime},\omega^{\prime})\in{\mathcal{H}}_{g-1}(2^{g-2},0,0)$. if we want to collapse any $k$ closed saddle connections, we have to ensure the locus of surgery for these saddle connections are disjoint. For a given interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, define the exception set $\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ to be the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ such that on every $(X,\omega)\in\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ there are two closed saddle connections with length in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ whose surgery loci intersect. First we will show that the measure of exception set goes to zero as genus goes to infinity.
5.1. The measure of exception set
For $(X,\omega)\in\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, there are three situations:
  • 1
    There exist two closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ that do not share a zero. And the loci of opening surgery intersect or the two closed saddle connections intersect.
  • 2
    There exist two closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ sharing one zero.
Note that in the first situation, we can find a curve connecting the two zeros with length no more than $\frac{2b}{\sqrt{g}}$.
Fix $B\in\mathbb{R}^{+}$, define $\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})$ be the set of $(X,\omega)\in\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[0,\frac{B}{\sqrt{g}}]$ such that there exist two closed saddle connections on $(X,\omega)$ that do not share a zero, moreover the lengths of the two closed saddle connections and the distance between the two zeros are less than $\frac{B}{\sqrt{g}}$. Define $\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}})$ be the set of $(X,\omega)\in\mathcal{C}^{1,0}_{g,Per}(2^{g-1})[0,\frac{B}{\sqrt{g}}]$ such that there exists two closed saddle connections with lengths less than $\frac{B}{\sqrt{g}}$ sharing one zero. Obviously for $B\geq 2b$, we have
$\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]\subset\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}).$
We need to compute the measure of $\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})$ and $\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}})$ when $g\to\infty$. First consider $\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}})$, we have